GreenHab Report – December 10th

Greenhab Officer:
Makiah Eustice

Environmental Control:
Heating
40% Shade Cloth on

Average Temperatures:
Low: 24 C
High: 47 C

Hours of Supplemental Light:
5

Daily Water Usage of Crops:
6 gallons

Water in Blue Tank (lbs):
70%

Times of Watering for Crops:
1430

Changes to crops:
NONE

Narrative:
Plants look happier with more water.

Harvest:
2.4 g of Basil
23.6 g of Mixed salad greens.

Support/Supplies Needed:
Any ideas for sensors to put in the water tank to know the level? The water gnats are getting out of hand. The sticky sheets are not effective enough.

Journalist Report – December 10th

Crew 200 Journalist Report 10-DEC-2018

On the last day of our nominal mission, we filmed a video that introduced the habitat and its surrounding buildings. We filmed it using Italian, Chinese, French, Japanese, German, and English, with each crew member hosting a part of it. We had to “rehearse” it beforehand to film it in one continuous shot. It was a memorable way to finish our mission, highlight the diverse nature of our team, and illuminate that MDRS is for the global public. We strive to continue demonstrating these values during our extended mission.

Today marks the first simulation date for our reduced three-member crew. Roles that four members impeccably served during the nominal mission are now vacant. We need to be fast-learning jacks and jills of all trades, even more so now than we did before. With the sudden departure of crew members who were culinarily inclined, something as simple as cooking could prove dismal. Thankfully, we have not resorted to eating habitat mice! We tap into skills acquired from the full crew and our creative juices to concoct delicious meals derived from dehydrated ingredients.

We were given the special opportunity to plan and perform the first EVA with a three-member crew, which we achieved today. We applied a communication protocol where we recorded QSA and QRK codes every minute and GPS coordinates every four minutes. Throughout the week, we aim to map the strength and readability of radio communication for areas surrounding the habitat. We will also develop a risk analysis matrix for three-member EVAs based on our experiences. We are motivated to contribute fresh knowledge in the area of reduced crew EVAs for the benefit of MDRS and beyond.

Tonight we celebrate our successful EVA with spaghetti and pesto, garnished with basil leaves from the GreenHab. We will also harvest the GreenHab to make a small and hearty salad. Admittedly sprinkled with water gnats, also courtesy of the GreenHab. Thank you to CapCom and others back on Earth for supporting us on this three-person mission. We will not let you down!

EVA Report – December 10th

Crew 200 EVA Report 10-12-2018

EVA #3

Author: Antoine Bocquier

Purpose of EVA: 1) Map out locations around the habitat that allow for radio communication between EVA crew and HabCom, 2) Test how EVAs can safely and effectively be conducted with three-member crews.

Start time: 10:35

End time: 12:30

Narrative: Crew conducted first 3-crew member EVA to explore and map the regions of communication strength/readibility. We applied a
communication protocol to control the radio strength/readability (QSA/QRK codes) every minute and communicate GPS coordinates every 5mn. We planned contingency plans for loss of signal and navigation. Afterwards, the crew mapped the locations of perceived path by each crew member against the estimated path based on recorded GPS coordinates. We could not explore as much area as planned (we planned to have a walk to Phobos’s Peak) given the time left but the experiment was a success, both on the EVA crew and HabCom sides for a 3-member crew.

Destination: Kissing Camel East Ridge

Coordinates (use UTM NAD27 CONUS): 519000 E – 4249500 N to 518500 E – 42505000 N

Participants: Makiah, Lindsay

Road(s) and routes per MDRS Map : N/A

Mode of travel: walking

Operations Report – December 10th

Crew 200 Operations Report 10-Dec-2018

SOL: 8

Name of person filing report: Antoine Bocquier

Non-nominal systems: generator autostart
Notes on non-nominal systems: None

Generator (hours run): From 5:15 pm last night to 9:00 am. On tonight at 4:50 pm.

Solar— SOC 96%
Diesel Reading – 24 %
Propane Reading – 34 %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 380 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – refilled, high level
Water (loft) – Static to Loft Pump used – yes
Water Meter: 4

Toilet tank emptied: no

Deimos rover used: no, still not functional
Hours:141.5
Beginning charge:
Ending charge: 71%
Currently charging: no

Sojourner rover used: No
Hours:
Beginning charge:
Ending charge: 100%
Currently charging: yes

Spirit rover used: no
Hours: 59.6
Beginning charge:
Ending charge: 100%
Currently charging: yes

Opportunity rover used: no
Hours: 41.1
Beginning charge:
Ending charge: 100%
Currently charging: yes

Curiosity rover used: no
Hours: 57.5
Beginning charge:
Ending charge: 100%
Currently charging: yes

Notes on rovers: Rovers were not used this week
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station

HabCar used and why, where? No
CrewCar used and why, where? No

General notes and comments: Eveything looks fine

Summary of internet: Nothing to report.

Summary of suits and radios: we performed an EVA this morning to study the radio coverage close to the Kissing Camel range. Spacesuit 3 does not seem to work, I am trying another charger

Summary of Hab operations: everything is fine, living, cooking and sleeping are still pleasant.

Summary of GreenHab operations: Regular activities

Summary of ScienceDome operations: We mopped the ground.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Musk Solar observatory – Nominal. Robotic Observatory – Was open all night and day long, we exchanged with Peter and CapCom to fix this and it finally closed around 5:00pm.

Summary of health and safety issues: security checks performed

Questions, concerns and requests to Mission Support: Nothing to report

Sol Summary – December 10th

Sol:8

Summary Title: The first 3 crew EVA on Mars

Author’s name: Makiah Eustice

Mission Status: Crew conducted an EVA with 2 crew members out and only one member on base, mopped the Science Dome.

Sol Activity Summary: Crew conducted first 3-crew member EVA to explore and map the regions of communication strength. We planned contingency plans for loss of signal and navigation. Afterwards, the crew mapped the locations of perceived path by each crew member against the estimated path based on recorded GPS coordinates. The rest of the day was relaxed and quiet, taking time to clean the mudded floor of the Science Dome and have computer-based work on our projects.

Look Ahead Plan: Tomorrow an driving/walking EVA is planned. Crew projects will continue collaborative projects.

Anomalies in work: The observatory was open most of the day. It closed once we restarted the generator.

Weather: Cold, muddy from snow melt.

Crew Physical Status: No crew members reported physical problems today.

EVA: EVA crew went, by foot, to Kissing Camel Ridge East to explore the limits of walkie talkie comms in certain regions. Comms checks were performed every minute between HabCom and EVA, and GPS coordinate was exhanged every few minutes. EVA was planned for 2 hours and was conducted in exactly 2 hours.

Reports to be file: Operations Report, EVA Report, EVA Request, GreenHab Report, Pictures Report, Journalist Report, HSO Report

Support Requested: N/A